There are questions and answers. I want to know explanation how to get this answers. Thant you:)
This is it;
[Situation]
-The home country is currently in its LR eq'm with:
P=1, M1s=1, G1=T1=0, Y1=Ypotential
-Foreign variables: i*=2(%), P*=1
-Assume that the IS curve is a straight line, and that its slope is preserved.
-MPC=0.75
-Investment depends negatively on i.
-TB=-0.25xY+0.5x(EP*/P)
-Money demand: Md=Px(Y-i)
-The initial IS curve labelled IS1 is shown in the handout.
-In the LR:
1) PPP holds, and
2) the expected ER is equal to the spot ER.

Q. Given the initial values of P=1 and M1s=1, find the equation of the initial LM1 curve.
A. i=Y-1

Q. Using the initial IS1 curve in the handout and the LM1 curve obtained in Q1, find the values of (Y,i) in the initial eq'm .
A. (3,2)

Q. Given the "Equilibrium1", find the values of (i, E, Ee) in the corresponding initial eq'm in the FX market.
A. (2,1,1)
1
Expert's answer
2016-12-16T06:18:11-0500
All the answers were obtained according to the data quoted in this task because, for example, i = Y-1, we have a formula Md = Px (Yi), Md and Px is 1, and here we have what others The following answers are as specified in the job.